# Generating entanglement of two acoustic modes by driving the qubit in circuit quantum acoustodynamics system

### v1

@@ II Theoretical model and the effective Hamiltonian
Figure 1: (a) Schematic representation of a circuit quantum acoustodynamics system. The HBAR chip, positioned at the top, is equipped with a layer of piezoelectric aluminum nitride (green) and supports two acoustic modes (blue and red). The superconducting transmon qubit on the lower chip couple with the HBAR. (b) Frequency spectrum of the system. The qubit with frequency $\omega_{q}$ is driven by a laser $\omega_{d}$, where $\delta_{d}=\omega_{q}-\omega_{d}$.
@@ III The quantitative results of the acoustic-mode entanglement
Figure 5: (a) Density plot of the maximum entanglement $E_{N}$ versus the phonon dissipation $\gamma$ and temperature $T$. (b) The maximum entanglement $E_{N}$ versus the temperature $T$ for different values of $\kappa_{q}$. We take $\Omega_{d}=25$ $\rm MHz$ and the other parameters are the same as those in Fig. 2.
$E_{N}=\mathrm{max}[0,-\mathrm{ln}(2\eta^{-})],$ (17)
@@
Now we discuss the time evolution of entanglement between two bosonic modes by numerically solving the master equation with the full Hamiltonian $(H_{3})$ and the effective Hamiltonian $(H_{7})$, respectively. The experimentally feasible parameters are selected 29 : $\omega_{a}/2\pi=5.9236$ $\rm GHz$, $\omega_{b}/2\pi=5.9488$ $\rm GHz$, $g=g_{a}=g_{b}=257$ $\rm kHz$, $\gamma_{a}/2\pi=4.7$ $\rm kHz$, and $\gamma_{b}/2\pi=3.3$ $\rm kHz$. These parameters can also ensure the system is operated in the phonon-qubit dispersive coupling regime $g\ll|\omega_{q}-\omega_{a,b}|$. Here we temporarily assume that $\kappa_{q}=0$ and $T=50$ $\rm mK$. Initially, the phonon modes are in the vacuum states, and the qubit is prepared in the state |+ket|$\rangle$| + ⟩.
@@
The evolutions of entanglement over time $t$ for different driving amplitudes $\Omega_{d}$ are plotted in Fig. 2(a), which shows that the considerable entanglement between two acoustic modes can be achieved after evolving for tens of microseconds. Moreover, the results obtained using the full Hamiltonian (solid and dash-dot lines) and the effective Hamiltonian (dashed and dotted lines) agree well with each other, indicating that the approximation used for deriving the effective Hamiltonian is very reasonable. Meanwhile, one can see the entanglement for $\Omega_{d}=25$ MHz is greater than that for $\Omega_{d}=15$ MHz, which can be explained through Eq. (10). In Eq. (10), the effective coupling strength $G=\frac{\Omega_{d}g^{2}}{\delta^{2}-4\Omega_{d}^{2}}$. Obviously, $\Omega_{d}$ cannot equal to $|\delta|/2$ which causes the system to infinity, and the effective coupling strength will increase with increasing $\Omega_{d}$ for $\Omega_{d}<|\delta|/2$, leading to the increase of the entanglement. Nevertheless, the driving strength $\Omega_{d}$ and detuning $\delta_{d}$ cannot be arbitrarily chosen, because they are also constrained by Eq. (9). To discuss the effect of $\Omega_{d}$ and $\delta_{d}$ on the entanglement, we plot the entanglement as functions of the detuning $\delta_{d}$ for different $\Omega_{d}$ and $t=100$ $\mu s$ in Fig. 2(b). One can see, for a given $\Omega_{d}$, there is an optimal $\delta_{d}$, which corresponds to Eq. (9) being satisfied. On the other hand, the entanglement can be present for a wide range of $\delta_{d}$, which means that the requirements for experimental parameters in this scheme are relatively relaxed, i.e., Eq. (9) does not need to be strictly satisfied.
@@
What should be emphasized is that, to demonstrate the validity of the effective Hamiltonian, we have chosen the decay rate of qubit $\kappa_{q}=0$ in Fig. 2. In practice, $\kappa_{q}$ is bound to affect the performance of the scheme, so it must be taken into account. In Fig. 3(a), we plot the entanglement dynamics for different $\kappa_{q}$ by using the full Hamiltonian (the results obtained using the effective Hamiltonian is not affected by $\kappa_{q}$). From Fig. 3(a), we can see the obtained entanglement will decay to 0 with increasing the interaction time, and the smaller $\kappa_{q}$, the stronger entanglement that can be obtained. That’s because the acoustic modes will dissipate through the qubit due to the exchange interaction between them, although the dissipation rate of the acoustic mode is very small. Therefore, to make the entanglement continue for a long time, the interaction between acoustic modes and the qubit can be turned off when the entanglement reaches the maximum. In order to visually show the impact of system dissipations, we choose the two acoustic modes has the same dissipation rate $\gamma=\gamma_{a}=\gamma_{b}$, and plot the maximum entanglement versus $\kappa_{q}$ and $\gamma$ in Fig. 3(b), which shows the degree of entanglement is significantly affected by $\kappa_{q}$. Fortunately, the qubit decay rate of $\kappa_{q}/2\pi\sim 10$ kHZ and the phonon dissipation of $3~{}4$ kHZ have been experimentally achieved in cQAD systems 29 ; 30 . Therefore, Fig. 3(b) indicates the entanglement can be obtained within the range of experimentally feasible parameters, even if $\kappa_{q}$ reaches $2\pi\times 30$ kHZ.

### v2

@@ II Theoretical model and the effective Hamiltonian
Figure 1: (a) Schematic of the circuit quantum acoustodynamics system. The HBAR chip, positioned at the top, is equipped with a layer of piezoelectric aluminum nitride (green) and supports two acoustic modes (blue and red). The superconducting transmon qubit on the lower chip couple to the HBAR through piezoelectric interaction. (b) Frequency spectrum of the system. The qubit with frequency $\omega_{q}$ is driven by a microwave field with frequency $\omega_{d}$, where $\delta_{d}=\omega_{q}-\omega_{d}$.
@@ III The quantitative results of the acoustic-mode entanglement
Figure 5: (a) Density plot of the maximum entanglement $E_{N}$ versus the phonon dissipation $\gamma$ and temperature $T$ for $\kappa_{q}/2\pi=10$ $\rm kHz$. (b) The maximum entanglement $E_{N}$ versus the temperature $T$ for different values of $\kappa_{q}$. We take $\Omega_{d}=25$ $\rm MHz$ and the other parameters are the same as those in Fig. 2.
@@
Now we discuss the time evolution of entanglement between two bosonic modes by numerically solving the master equation with the full Hamiltonian $(H_{3})$ and the effective Hamiltonian $(H_{7})$, respectively. The experimentally feasible parameters are selected 29 : $\omega_{a}/2\pi=5.9236$ $\rm GHz$, $\omega_{b}/2\pi=5.9488$ $\rm GHz$, $g/2\pi=g_{a}/2\pi=g_{b}/2\pi=257$ $\rm kHz$, $\gamma_{a}/2\pi=4.7$ $\rm kHz$, and $\gamma_{b}/2\pi=3.3$ $\rm kHz$. These parameters can also ensure the system is operated in the phonon-qubit dispersive coupling regime $g\ll|\omega_{q}-\omega_{a,b}|$. Here we temporarily assume that $\kappa_{q}=0$ and $T=50$ $\rm mK$. Initially, the phonon modes are in the vacuum states, and the qubit is prepared in the state |+ket|$\rangle$| + ⟩.
@@
The evolutions of entanglement over time $t$ for different driving amplitudes $\Omega_{d}$ are plotted in Fig. 2(a), which shows that the considerable entanglement between two acoustic modes can be achieved after evolving for tens of microseconds. Moreover, the results obtained using the full Hamiltonian (solid and dash-dot lines) and the effective Hamiltonian (dashed and dotted lines) agree well with each other, indicating that the approximation used for deriving the effective Hamiltonian is very reasonable. Meanwhile, one can see the entanglement for $\Omega_{d}=25$ MHz is greater than that for $\Omega_{d}=15$ MHz, which can be explained through Eq. (10). In Eq. (10), the effective coupling strength is $G=\frac{\Omega_{d}g^{2}}{\delta^{2}-4\Omega_{d}^{2}}$. Obviously, $\Omega_{d}$ cannot equal to $|\delta|/2$ which causes the system to infinity, and the effective coupling strength will increase with increasing $\Omega_{d}$ for $\Omega_{d}<|\delta|/2$, leading to the increase of the entanglement. Nevertheless, the driving strength $\Omega_{d}$ and detuning $\delta_{d}$ cannot be arbitrarily chosen, because they are also constrained by Eq. (9). To discuss the effect of $\Omega_{d}$ and $\delta_{d}$ on the entanglement, we plot the entanglement as functions of the detuning $\delta_{d}$ for different $\Omega_{d}$ and $t=100$ $\mu s$ in Fig. 2(b). One can see, for a given $\Omega_{d}$, there is an optimal $\delta_{d}$, which corresponds to Eq. (9) being satisfied. On the other hand, the entanglement can be present for a wide range of $\delta_{d}$, which means that the requirements for experimental parameters in this scheme are relatively relaxed, i.e., Eq. (9) does not need to be strictly satisfied.
@@
What should be emphasized is that, to demonstrate the validity of the effective Hamiltonian, we have chosen the decay rate of qubit $\kappa_{q}=0$ in Fig. 2. In practice, $\kappa_{q}$ is bound to affect the performance of the scheme, so it must be taken into account. In Fig. 3(a), we plot the entanglement dynamics for different $\kappa_{q}$ by using the full Hamiltonian (the results obtained using the effective Hamiltonian is not affected by $\kappa_{q}$). From Fig. 3(a), we can see the obtained entanglement will decay to 0 with increasing the interaction time, and the smaller $\kappa_{q}$, the stronger entanglement that can be obtained. That’s because the acoustic modes will dissipate through the qubit due to the exchange interaction between them, although the dissipation rate of the acoustic mode is very small. Therefore, to make the entanglement continue for a long time, the interaction between acoustic modes and the qubit can be turned off when the entanglement reaches the maximum. In order to visually show the impact of system dissipations, we choose the two acoustic modes has the same dissipation rate $\gamma=\gamma_{a}=\gamma_{b}$, and plot the maximum entanglement versus $\kappa_{q}$ and $\gamma$ in Fig. 3(b), which shows the degree of entanglement is significantly affected by $\kappa_{q}$. Fortunately, the qubit decay rate of $\kappa_{q}/2\pi\sim 10$ kHz and the phonon dissipation of $3\sim 4$ kHz have been experimentally achieved in cQAD systems 29 ; 30 . Therefore, Fig. 3(b) indicates the entanglement can be obtained within the range of experimentally feasible parameters, even if $\kappa_{q}$ reaches $2\pi\times 30$ kHz.